Solve the system by graphing:
y=5x-6
y=-4x+3

Answer by mananth(16946)

y=5x-6
y=-4x+3
equation1 -6 = -6
when x= 0 y= -6 ( 0 , -6 )
when y = 0 x= 1.2 ( 1.2 , 0)
Equation 2
when x= 0 y= 3 ( 0 , 3 ) 3
when y = 0 x= 0.75 ( 0.75 , 0 )